OpenStudy (purplemexican):

Which values of p and q would make the value of the following expression equal to 29i? (2 – 5i)(p + q)i

OpenStudy (solomonzelman):

Ok, expand the (2-5i)(p+q)i and what do you then get?

OpenStudy (purplemexican):

expand?

OpenStudy (lynfran):

yes in other words multiply out the bracket

OpenStudy (purplemexican):

its still confusing

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle (a+b)(c+d)=ab+ac+bc+bd }\) have you seen anything like this before?

OpenStudy (purplemexican):

yes

OpenStudy (purplemexican):

i need to look into which property is which still get them confused

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle (2 – 5i)(p + q)=2(p+q)-5i(p+q) }\) same way as: \((a-b)(C)=a(C)-b(C)\) ---------------------------------------- And therefore: \(\large\color{black}{ \displaystyle (2 – 5i)(p + q)i=2i(p+q)-5i^2(p+q) }\) we know \(i^2=-1\) because \(i^2=\left(\sqrt{-1}\right)^2=-1,\) like any \(\left(\sqrt{a}\right)^2=a\) So, \(\large\color{black}{ \displaystyle (2 – 5i)(p + q)i=2i(p+q)-5(-1)(p+q) }\) \(\large\color{black}{ \displaystyle (2 – 5i)(p + q)i=2i(p+q)+5(p+q) }\)

OpenStudy (solomonzelman):

Now, how would you get 29i in a "a+bi" form? 0+29i and so here we have this form as well 5(p+q) + 2(p+q)i so a=0=5(p+q) bi=29i=2(p+q)i

OpenStudy (solomonzelman):

solve the system 0=5(p+q) 29=2(p+q) (---> this eq. results after canceling the i's on both sides.)

OpenStudy (solomonzelman):

good luck